# HeadEvolver: Text to Head Avatars via Expressive and Attribute-Preserving Mesh Deformation

### v2

@@ Abstract
Current text-to-avatar methods often rely on implicit representations, leading to 3D content that artists cannot easily edit and animate in graphics software. This paper introduces a novel framework for generating stylized head avatars from text guidance, which leverages locally learnable mesh deformation and 2D diffusion priors to achieve high-quality digital assets for attribute-preserving manipulation. Given a template mesh, our method represents mesh deformation with per-face Jacobians and adaptively modulates local deformation using a learnable vector field. This vector field enables anisotropic scaling while preserving the rotation of vertices, which can better express identity and geometric details. We also employ landmark- and contour-based regularization terms to balance the expressiveness and plausibility of generated head avatars from multiple views without relying on any specific shape prior. Our framework can not only generate realistic shapes and textures that can be further edited via text, but also support seamless editing using the preserved attributes from the template mesh, such as 3DMM parameters, blendshapes, and UV coordinates. Extensive experiments demonstrate that our framework can generate diverse and expressive head avatars with high-quality meshes that artists can easily manipulate in 3D graphics software, facilitating downstream applications such as more efficient asset creation and animation with preserved attributes. Our project page is at: https://www.duotun-wang.co.uk/HeadEvolver/
@@ 1 Introduction
However, there is a significant gap between current text-to-avatar methods and 3D asset creation, as most methods do not effectively integrate high fidelity, user customization, and animation support. This is mainly caused by using implicit shape representations that cannot preserve 3D attributes essential for downstream manipulation, e.g., mesh topology, rig, UV mapping. For instance, HeadArtist [31] and HeadSculpt [15] leverage DMTet [40] to create head avatars with 3D Morphable Models (3DMMs) [4, 27] as strong shape priors. While this approach allows for a realistic appearance, the implicit shape representation DMTet used in these works tends to discard statistical information such as shape and expression parameters from 3DMMs and produce noisy meshes.
@@
Head avatar generation using an explicit mesh representation would benefit downstream applications because of its modifiability and compatibility with graphics pipelines, although there is less existing research in this direction. A notable method is TADA [29], which incorporates the optimizations of vertex displacement and SMPL-X parameters [35]. This strategy preserves the semantics of 3DMMs and enables animatable 3D head avatar generation. However, direct vertex displacement tends to deform meshes with self-intersected faces and noisy normals, leading to limited mesh quality and animation artifacts. TextDeformer [10] can change an input mesh smoothly and stably by making deformation via Jacobians, but it cannot generate a textured appearance and lacks local geometry details.
@@
This paper aims to mitigate the trade-offs between creating high-quality avatars and the ease of animation and editing without reconstructing blendshapes or rigs. We propose a novel text-to-avatar framework that produces stylized head avatars through expressive and attribute-preserving mesh deformation. Given a single template mesh such as FLAME [27], ICT-Face [26], or any manifold mesh preferably with facial semantics, our framework deforms it into a desired target shape while preserving feature correspondences such as skinny weights and UV coordinates for further animation and editing (Fig. LABEL:fig:teaser(c)).
@@
The key observation is that although the mesh deformation via optimizing Jacobians supports smooth shape changes at a global scale [1], it falls short in fully expressing the semantics by optimizing image losses and performing locally anisotropic deformation. This is essential for generating diverse styles, such as the elongated nose in Fig. LABEL:fig:teaser(b). Furthermore, it is often laborious to tune deformation hyper-parameters to achieve both text-guided expressiveness and shape smoothness [34]. Therefore, we propose per-triangle learnable weighting factors coupled with Jacobians to enhance fine-grained deformation over the head mesh, which we call vector fields.
@@
In order to build a complete 3D head avatar generation framework via gradient-based optimization, we employ a pre-trained diffusion model [36] with 3D awareness by leveraging a frozen landmarkguided ControlNet [57, 33]. Given a camera pose, we render shaded and normal images and obtain the opacity mask and the predefined vertex-indexed landmarks projected from the 3D head. As Fig. 1 illustrates, the stable diffusion model accepts text prompts, rendered images, and landmarks as additional conditions from ControlNet to compute the Score Distillation Sampling (SDS) loss. In addition, to mitigate the issues of unexpected geometry distortions or self-intersected triangles, we introduce landmark- and contour-based regularization terms to properly constrain the avatar deformation. For example, a good-shaped avatar is likely to be vertically symmetric. To summarize, we make the following contributions:
@@
  • We introduce per-triangle vector fields to enhance the expressiveness of mesh deformation over local regions while preserving 3D attributes.
@@
  • We propose a diffusion-based framework with avatar regularization terms to support text-to-avatar generation with high-quality mesh.
@@ 2 Related Work
Table 1: Compared to other text-to-avatar approaches, our method preserves predefined 3D attributes (e.g., vertex-indexed landmarks, rig, blendshapes), is independent of parametric models for joint optimization, produces mesh models of head avatars compatible with further manipulation in 3D software, and requires no training data nor fine-tuning.
3D Attribute Independence from Compatible with No Training Data
Preservation Parametric Models Graphics Pipelines nor Fine-tuning
@@
Recently, there has been rapid research progress in digital human modeling and text-to-3D content generation. Our research centers on text-guided synthesis of head avatars employing an explicit mesh representation, aiming to fulfill four criteria outlined in Table 1: 1) preservation of predefined 3D attributes, 2) support for template mesh independent from parametric models, 3) seamless compatibility with graphics pipelines, and 4) gradient-based optimization without requiring training data nor fine-tuning. This section reviews related literature on text-to-3D avatar generation and learning-based mesh optimization.
@@ 2.1 Text-to-3D Avatar Generation
The success of text-driven generative models has sparked a wide range of research works in 3D content synthesis. These methods are based on different geometry representations such as NeRF [30, 45], tri-plane [41], mesh[6], and Gaussian Splatting [43, 28]. For text-to-avatar generation tasks, numerous efforts have been made to develop 3D generative models [44, 56, 51, 53] typically based on GAN [13] or diffusion model [16]. These methods often demand extensive 3D head data for training, incurring high costs. Moreover, they face challenges due to limited 3D datasets for creating diverse head avatars matching text prompts.
To bypass the extensive training process of 3D generative models and insufficient 3D data, recent studies [18, 55] have made remarkable advances by harnessing the capabilities of vision-language models (e.g., CLIP [37] or SDS [36]) to generate static human avatars from text prompts. DreamWaltz [21] aims to enhance the realism of avatar animations with pose-guided ControlNet [57]. Similarly, DreamHuman [25] employs a signed distance field conditioned on pose and shape parameters to empower animations with NeRF. However, extracting, editing, and animating an explicit mesh from these works are not straightforward [21] (e.g., use marching cube or tetrahedra algorithms) since NeRF and DMTet [40], for instance, represent shapes through network weights. Meanwhile, the issue of poor mesh quality is often concealed by texture, as the geometry and texture optimizations are fully mixed. Recent works [6, 15, 31] disentangle the learning of geometry and texture to improve the avatar mesh quality. Our work also leverages differentiable rendering, stable diffusion, and a hybrid optimization process for mesh and texture synthesis, specifically emphasizing the deformation of an explicit geometry template through semantic guidance instead of generation from scratch or using implicit representations.
@@ 2.2 Learning-Based Mesh Optimization
Parametric mesh templates have been a prominent research focus for human avatar reconstruction and generation tasks, as they are well-suited for deep learning approaches. AvatarClip [19] and CLIP-Face [2] employ parametric models such as SMPL-X [35] and FLAME [27] as geometry prior to produce text-aligned avatars. DreamAvatar [5] extracts the shape parameters of SMPL as a 3D prior to learn a NeRF-based color field. HeadSculpt [15] and HeadArtist [31] utilize the FLAME template as canonical shape input and extract its facial landmark to guide avatar generations via DMTet. However, meshes extracted from implicit representations tend to lose semantic information of 3DMMs [4], posing challenges for downstream applications such as interactive editing and animation.
For manipulating geometry through deformation, recent works propose data-driven approaches [3, 1] to predict realistic deformation and some methods focus on learning deformation over one template mesh through tuning per-vertex coordinates [50, 58, 9]. TADA [29] directly optimizes vertex displacements along with the SMPL-X’s pose and shape parameters and achieves animatable avatars. Nonetheless, this characteristic is highly dependent on the quality of 3DMMs. Straightly optimization of non-convex mesh structures often converges to undesirable local minima, resulting in noticeable artifacts on surface details. TextDeformer [10] initially integrates CLIP into a learning-based deformation pipeline to address the scarcity of 3D datasets that pair shapes with text semantics. Our approach extends this by optimizing local semantic deformations on any template mesh, independent of its shape, expression, or blendshapes [4]. Our goal is to enhance identity expressiveness guided by text. The resulting head avatars can be animated using the original rigging from the source mesh and seamlessly integrated into existing rendering and editing workflows for design and artistic purposes.
@@ 3 Methodology
In this section, we describe the details of our generation pipeline, as shown in Fig. 1. Given a source template mesh, we jointly optimize the mesh deformation and an albedo map guided by a text prompt. Our approach is not limited to a specific parametric model (e.g., FLAME) and can be applied to any manifold mesh. Figs. 15 and 16 show the generalization of our approach.
@@
Figure 3: Qualitative comparisons. Our method preserves the template mesh’s 3D attributes, yielding diverse and high-quality mesh models.
3.1 Mesh Deformation through Jacobian Fields
Let $\mathcal{M}=(\mathcal{V},\mathcal{F})$ denote a source template mesh which consists of a set of vertices $\mathcal{V}\in\mathbb{R}^{n\times 3}$ and faces $\mathcal{F}\in\mathbb{Z}^{m\times 3}$. For the optimization strategy of mesh deformation, directly applying displacement to each vertex can result in severe self-interactions of triangle faces since it is susceptible to localized noisy gradients from pixel-level losses. Instead, we parameterize deformation mapping $\Phi:\mathbb{R}^{n\times 3}\to\mathbb{R}^{n\times 3}$ by employing a Jacobian field [1] $J=\{J_{f}|f\in\mathcal{F}\}$ that represents the scaling and rotation of each mesh face. A dual representation of $\mathcal{M}$ can be defined as a stack of per-face Jacobians $J_{f}\in\mathbb{R}^{3\times 3}$ where
@@ 3.1 Mesh Deformation through Jacobian Fields
and $\nabla_{f}$ is the gradient operator of triangle $f$. Conversely, we aim to optimize each vertex’s position from the per-face Jacobian.
@@
where $\nabla_{f}(\Phi)$ is the Jacobian of $\Phi$ at a triangle face $f$ and $|f|$ is the area of the face. Thus, the deformation map $\Phi$ over the input template mesh is indirectly optimized through Jacobians $J_{f}$ in the least square sense. The solution can be obtained by using a differentiable solver [1, 10] implementing Cholesky decomposition:
@@
where $\nabla$ is a stack of per-face gradient operators, $\mathcal{A}\in\mathbb{R}^{3m\times 3m}$ is the mass matrix of $\mathcal{M}$, and $L\in\mathbb{R}^{3n\times 3n}$ is the cotangent Laplacian of $\mathcal{M}$. This learnable shape representation based on Jacobians is robust to noisy gradients. It can achieve better geometry quality (e.g., fewer self-intersections and noisy normals) than direct vertex displacements, which is further justified in Sec. 4.
However, during our experiment, we observed that the poisson solver applied in Eq. 2 can impose strong geometry constraints for Jacobians to reach the global-coherence mesh deformation [10], which may lose fine-grain details for local shape features and compromise the identity of the generated head mesh (Fig. 5). A straightforward approach is to increase the learning rate for Jacobians to enhance their sensitivities to propagated gradients. This amplifies the rotation and scaling of each face, resulting in stronger deformations (Sec. 4.3). However, intensified rotations raise the likelihood of inverted triangles [39], leading to an unstable optimization process and a poor-quality mesh. To address these concerns, we introduce a learnable per-face vector field $H=\{H_{f}|f\in\mathcal{F}\}$ to enhance the deformation expressiveness of Jacobian field $J$, i.e.,
@@
By substituting $H$ into Eq. 3, the deformation mapping could be computed by
@@
The effect of $W=\{W_{f}|f\in\mathcal{F}\}$ is to control the anisotropic scaling (e.g., stretching and shrinking in different directions) on each face. As shown in Fig. 2, it can adaptively learn the degree of anisotropic scaling on the corresponding triangular faces that can represent the character’s features. Thus, the per-face vector field $H$ can prioritize the Poisson solver to approach the desired deformations to express the identity of head avatars more effectively than applying Jacobians only. The detailed validation is in Sec. 4.3.
@@ 3.2 Text-Guided 3D Synthesis
We aim to build a text-guided head avatar generation framework based on differentiable mesh deformation and texture generation. Motivated by recent success in large-scale VLMs and text-to-3D literature discussed in Sec. 2.1, we utilize a powerful 2D diffusion prior (i.e., SDS [36]) to direct deformations by scoring the rendering realism of the generated shape in our framework.
Specifically, by following the procedure described in Sec. 3.1, we denote $\mathcal{V^{*}}$ as the set of deformed vertices from $J$. Then, by using a differentiable render $\mathbf{R}$, we render $\mathcal{M^{*}}=(\mathcal{V}^{*},\mathcal{F})$ from a viewpoint delineated by camera extrinsic parameter $C$, which produces an image $I=\mathbf{R}(\mathcal{M^{*}},C)$. The SDS loss is leveraged to alter the geometry and color respectively, that is:
$\footnotesize\nabla_{J}\mathcal{L}_{\mathrm{SDS}}(\epsilon,I)=\mathbb{E}_{t,\epsilon}\left[w(t)((\epsilon_{\theta}(z_{t};y,C,t)-\epsilon)\frac{\partial{I}}{\partial{J}}\right],$ (6)
where $\epsilon\sim\mathcal{N}(0,I)$, $\epsilon_{\theta}$ is the pre-trained diffusion priors with the network parameter $\theta$, $w(t)$ is a weight coefficient related to timestep $t\sim\mathcal{U}(0,1)$, $y$ is the text input condition, $z_{t}$ denotes a latent embedding of $I$ perturbed with noises at time step $t$, and $C$ is the MediaPipe [33] facial landmark map. Rendered image $I$ can be obtained from the normal $I_{n}$ and shaded $I_{s}$ (Fig. 1).
To ensure deformation mapping from propagated gradients and the Poisson solver can conform to the facial topology and reasonable semantics of the source mesh, we further develop two regularization terms, aiming to preserve 3D attributes of generated head avatars (Figs. 12 and 13).
@@
$\mathcal{L}=\lambda_{1}\nabla_{\Phi}\mathcal{L}_{\mathrm{SDS}}+\lambda_{2}\mathcal{L}_{lmk}+\lambda_{3}\mathcal{L}_{op},\vspace{-0.05cm}$ (7)

### v3

@@ Abstract
Current text-to-avatar methods often rely on implicit representations (e.g., NeRF, SDF, and DMTet), leading to 3D content that artists cannot easily edit and animate in graphics software. This paper introduces a novel framework for generating stylized head avatars from text guidance, which leverages locally learnable mesh deformation and 2D diffusion priors to achieve high-quality digital assets for attribute-preserving manipulation. Given a template mesh, our method represents mesh deformation with per-face Jacobians and adaptively modulates local deformation using a learnable vector field. This vector field enables anisotropic scaling while preserving the rotation of vertices, which can better express identity and geometric details. We employ landmark- and contour-based regularization terms to balance the expressiveness and plausibility of generated avatars from multiple views without relying on any specific shape prior. Our framework can generate realistic shapes and textures that can be further edited via text, while supporting seamless editing using the preserved attributes from the template mesh, such as 3DMM parameters, blendshapes, and UV coordinates. Extensive experiments demonstrate that our framework can generate diverse and expressive head avatars with high-quality meshes that artists can easily manipulate in graphics software, facilitating downstream applications such as efficient asset creation and animation with preserved attributes. Our project page is at: https://www.duotun-wang.co.uk/HeadEvolver/
@@ 1 Introduction
However, there is a significant gap between current text-to-avatar methods and 3D asset creation, as most methods do not effectively integrate high fidelity, user customization, and animation support. This is mainly caused by employing implicit shape representations that cannot preserve 3D attributes essential for downstream manipulation, e.g., mesh topology, rig, and UV mapping. For instance, HeadArtist [32] and HeadSculpt [14] leverage DMTet [44] to create head avatars with 3D Morphable Models (3DMMs) [4, 28] as shape priors. While these approaches create a realistic appearance, DMTet discards statistical parameters from 3DMMs, leading to noisy meshes.
@@
Head avatar generation using an explicit mesh representation would benefit downstream applications because of its editability and compatibility with graphics pipelines, although there is less existing research in this direction. A notable method is TADA [30], which incorporates the optimizations of vertex displacement and SMPL-X parameters [36]. This strategy preserves the semantics of 3DMMs and enables animatable 3D head avatar generation. However, direct vertex displacement tends to deform meshes with self-intersected faces and noisy normals, leading to limited mesh quality and artifacts in the animation. TextDeformer [9] can change an input mesh smoothly and stably by deforming through Jacobians, but it cannot generate a textured appearance and lacks local geometry details.
@@
This paper aims to create high-quality avatars that support animation and editing without reconstructing blendshapes or rigs. We propose a novel text-to-avatar framework that produces stylized head avatars through expressive and attribute-preserving mesh deformation. Given a single template mesh such as FLAME [28], ICT-Face [27], or any manifold mesh preferably with facial semantics, our framework deforms it into a desired target shape while preserving feature correspondences such as skinny weights and UV coordinates for animation and editing, as shown in Figure LABEL:fig:teaser(c).
@@
The key observation is that optimizing Jacobians enables smooth global shape changes [1] but lacks expressive local deformations when processing gradients from text and image losses, compared to direct vertex displacements. This local control is essential for generating diverse styles, such as the elongated nose in Figure LABEL:fig:teaser(b). Furthermore, it is often laborious to tune deformation hyper-parameters to achieve both text-guided expressiveness and shape smoothness [35]. Therefore, we propose per-triangle learnable weighting factors coupled with Jacobians to enhance fine-grained deformations, which we refer to as vector fields. The notion of vector fields was initially introduced through a Poisson-based solver [57] to guide mesh deformations and reconstructions. In this study, we expand upon this by incorporating differentiable settings for 3D generation tasks.
@@
In order to build a complete 3D head avatar generation framework via gradient-based optimization, we employ a pre-trained diffusion model [38] with 3D awareness by leveraging a frozen landmark-guided ControlNet [63, 34]. Given a camera pose, we render shaded and normal images and obtain the opacity mask and the predefined vertex-indexed landmarks projected from the 3D head. As Figure 1 illustrates, the stable diffusion model accepts text prompts, rendered images, and landmarks as additional conditions from ControlNet to compute the Score Distillation Sampling (SDS) loss. In addition, to mitigate unexpected geometry distortions and self-intersecting triangles, we introduce landmark- and contour-based regularization terms. In summary, our contributions are as follows.
@@
  • We introduce a per-triangle vector field to enhance the expressiveness of mesh deformation over local regions while preserving 3D attributes.
@@
  • We propose a diffusion-based framework with landmark- and contour-based regularization terms for high-quality text-to-avatar mesh generation.
@@ 2 Related Work
Recently, there has been rapid research progress in digital human modeling and text-to-3D content generation. Our work centers on text-guided synthesis of head avatars employing an explicit mesh representation, aiming to fulfill three criteria outlined in Table 1. In addition, our framework generates mesh models that are compatible with graphic pipelines for direct manipulation in 3D software.
Table 1: Compared to other text-to-avatar approaches, our method preserves predefined 3D attributes (e.g., vertex-indexed landmarks, rig, blendshapes). It accommodates any template mesh independent of parametric models and operates solely on gradient-based optimizations without training data or fine-tuning.
Attribute Non-reliance on No Training Data
Preservation Parametric Models nor Fine-tuning
@@ 2.1 Text-to-3D Avatar Generation
The success of text-driven generative models has sparked a wide range of research in 3D content synthesis. These methods are based on different geometry representations such as NeRF [31, 49], tri-plane [45], mesh[6], and Gaussian Splatting [47, 29]. In text-to-avatar generation, many efforts have been made to develop 3D generative models [48, 62, 56, 59] typically based on GANs [12] or diffusion models [15]. These methods often demand extensive 3D head data for training, incurring high costs. Moreover, they face challenges due to limited 3D datasets for creating diverse head avatars matching text prompts.
To bypass the extensive training process of 3D generative models and insufficient 3D data, recent research [17, 61] has made remarkable advances by harnessing the capabilities of vision-language models (e.g., CLIP [41] and SDS [38]) to generate static human avatars from text prompts. DreamWaltz [21] aims to enhance the realism of avatar animations with pose-guided ControlNet [63]. Similarly, DreamHuman [26] employs a signed distance field conditioned on pose and shape parameters to empower animations with NeRF. However, extracting, editing, and animating an explicit mesh from these works are not straightforward [21] (e.g., use marching cube or tetrahedra algorithms) since NeRF and DMTet [44], for instance, represent shapes through network weights. Meanwhile, the issue of poor mesh quality is often concealed by texture, as the geometry and texture optimizations are fully mixed. Recent works [6, 14, 32] disentangle the learning of geometry and texture to improve the avatar mesh quality. Our work also leverages differentiable rendering, stable diffusion, and a hybrid optimization process for mesh and texture synthesis, specifically emphasizing the deformation of an explicit geometry template through semantic guidance instead of generation from scratch or using implicit representations.
@@ 2.2 Learning-Based Mesh Optimization
Parametric mesh templates have been a prominent research focus for human avatar reconstruction and generation tasks, as these models are well-suited in a deep learning context. AvatarClip [18] and CLIP-Face [2] employ parametric models such as SMPL-X [36] and FLAME [28] as geometry prior to produce text-aligned avatars. DreamAvatar [5] extracts the shape parameters of SMPL as a 3D prior to learn a NeRF-based color field. HeadSculpt [14] and HeadArtist [32] utilize the FLAME template as canonical shape input and extract its facial landmark to guide avatar generations through DMTet. However, meshes extracted from implicit representations tend to lose the semantics of 3DMMs [4], posing challenges for downstream applications such as interactive editing and animation.
From the perspective of geometry editing through explicit deformation, recent works propose data-driven approaches [3, 1] to predict realistic deformation and some methods focus on learning deformation over one template mesh through tuning per-vertex coordinates [54, 64, 8]. TADA [30] directly optimizes vertex displacements along with the SMPL-X’s pose and shape parameters and achieves animatable avatars. However, this characteristic is highly dependent on the quality of 3DMMs. Straightly optimization of non-convex mesh structures often converges to undesirable local minima, resulting in noticeable artifacts on surface details. TextDeformer [9] initially integrates CLIP into a learning-based deformation pipeline to address the scarcity of 3D datasets that pair shapes with captions. Our approach extends this by employing vector fields to optimize local semantic deformations on any template mesh, independent of its shape, expression, or blendshapes [4]. Our goal is to enhance identity expressiveness guided by text input. The resulting head avatars can be animated using the original rigging from the source mesh and seamlessly integrated into existing 3D design and animation workflows.
@@ 3 Methodology
In this section, we describe the details of our generation pipeline, as shown in Figure 1. Given a source template mesh, we jointly optimize the mesh deformation and an albedo map guided by a text prompt. Our approach is not limited to a specific parametric model (e.g., FLAME), and Figures 9 and 9 show the generalization of our approach.
@@
3.1 Mesh Deformation through Vector Fields
Let $\mathcal{M}=(\mathcal{V},\mathcal{F})$ denote a source template mesh which consists of a set of vertices $\mathcal{V}\in\mathbb{R}^{n\times 3}$ and faces $\mathcal{F}\in\mathbb{Z}^{m\times 3}$. For the optimization strategy of mesh deformation, directly applying displacement to each vertex can result in severe self-interactions of triangle faces since it is susceptible to localized noisy gradients from pixel-level losses. Instead, we parameterize deformation mapping $\Phi:\mathbb{R}^{n\times 3}\to\mathbb{R}^{n\times 3}$ by employing Jacobian fields [1] $J=\{J_{f}|f\in\mathcal{F}\}$ that represents the scaling and rotation of each mesh face. A dual representation of $\mathcal{M}$ can be defined as a stack of per-face Jacobians $J_{f}\in\mathbb{R}^{3\times 3}$ where
@@ 3.1 Mesh Deformation through Vector Fields
and $\nabla_{f}$ is the gradient operator of triangle $f$. Conversely, we optimize each vertex’s position from per-face Jacobian.
@@
where $\nabla_{f}(\Phi)$ is the Jacobian of $\Phi$ at a triangle face $f$ and $|f|$ is the area of the face. Thus, the deformation map $\Phi$ over the input template mesh is indirectly optimized through Jacobians $J_{f}$ in the least square sense. The solution can be obtained by using a differentiable solver [1, 9]:
@@
where $\nabla$ is a stack of per-face gradient operators, $\mathcal{A}\in\mathbb{R}^{3m\times 3m}$ is the mass matrix of $\mathcal{M}$, and $L\in\mathbb{R}^{3n\times 3n}$ is the cotangent Laplacian of $\mathcal{M}$. This learnable shape representation based on Jacobians is robust to noisy gradients. It can achieve better geometry quality (e.g., fewer self-intersections and noisy normals) than direct vertex displacements, which is further justified in Section 4.
However, during our experiment, we observed that the Poisson solver applied in Equation 2 can impose strong geometry constraints for Jacobians to reach the global-coherence mesh deformation [9], which may lose fine-grain details for local shape features and compromise the identity of the generated head mesh (Figure 5). A straightforward approach is to increase the learning rate for Jacobians to enhance their sensitivities to propagated gradients. This amplifies the rotation and scaling of each face, resulting in stronger deformations. However, intensified rotations raise the likelihood of inverted triangles [43], leading to an unstable optimization process and a poor-quality mesh. To address the above concerns, we introduce a learnable per-face vector field $H=\{H_{f}|f\in\mathcal{F}\}$ to enhance the deformation expressiveness of Jacobian field $J$, i.e.,
@@
By substituting $H$ into Equation 3, the deformation mapping could be computed by
@@
The effect of $W=\{W_{f}|f\in\mathcal{F}\}$ is to control the anisotropic scaling (e.g., stretching and shrinking in different directions) on each face. As shown in Figure 2, it can adaptively learn the anisotropic scaling on corresponding triangular faces that can represent the character’s features. Consequently, the per-face vector field $H$ prioritizes the Poisson solver for more effective head avatar identity expression than using only Jacobians (Section 4.4).
@@ 3.2 Text-Guided 3D Synthesis
We aim to build a text-guided head avatar generation framework based on differentiable mesh deformation and texture generation. Moti.vated by recent success in large-scale VLMs and text-to-3D literature discussed in Section 2.1, we utilize a powerful 2D diffusion prior (i.e., SDS [38]) to direct deformations by scoring the rendering realism of the generated shape in our framework.
Specifically, by following the procedure described in Section 3.1, we denote $\mathcal{V^{*}}$ as the set of deformed vertices from $J$. Then, by using a differentiable render $\mathbf{R}$, we render $\mathcal{M^{*}}=(\mathcal{V}^{*},\mathcal{F})$ from a viewpoint delineated by camera extrinsic parameter $C$, which produces an image $I=\mathbf{R}(\mathcal{M^{*}},C)$. The SDS loss is leveraged to alter the geometry and color respectively, that is:
$\footnotesize\nabla_{J}\mathcal{L}_{\mathrm{SDS}}(\epsilon,I)=\mathbb{E}_{t,\epsilon}\left[w(t)(\epsilon_{\theta}(z_{t};y,C,t)-\epsilon)\frac{\partial{I}}{\partial{J}}\right],$ (6)
where $\epsilon\sim\mathcal{N}(0,I)$, $\epsilon_{\theta}$ is the pre-trained diffusion priors with the network parameter $\theta$, $w(t)$ is a weight coefficient related to timestep $t\sim\mathcal{U}(0,1)$, $y$ is the text input condition, $z_{t}$ denotes a latent embedding of $I$ perturbed with noises at time step $t$, and $C$ is the MediaPipe [34] facial landmark map. Rendered image $I$ are obtained and encoded from the normal $I_{n}$ and shaded $I_{s}$, as seen in Figure 1.
To ensure deformation mapping from propagated gradients and the Poisson solver can conform to the facial topology and reasonable semantics of the source mesh, we further develop two regularization terms, aiming to preserve 3D attributes of generated head avatars (Figure 7).
@@
$\mathcal{L}=\lambda_{1}\nabla_{\Phi}\mathcal{L}_{\mathrm{SDS}}+\lambda_{2}\mathcal{L}_{lmk}+\lambda_{3}\mathcal{L}_{op},$ (7)